Two hoses fill a pool. 1st one fills the pool by itself in 4 hours/ The 2nd one fills it in 9 hours by itself.
How long does it take for both hoses to fill the pool?

Answer by richwmiller(17219)

This is a mixture word problem.
x/4+x/9=1
9x+4x=36
13x=36
x=36/13=2 10/13 =2 hrs and 46 minutes
36/13/4+36/13/9=1
9/13+4/13 =1